OpenStudy (anonymous):

Determine whether the equation defines y as a function of x. x^2 + y^2 = 1

OpenStudy (anonymous):

Can someone help me..

OpenStudy (anonymous):

\[ y^2 = 1 - x^2 \] Can you finish it now?

OpenStudy (anonymous):

so y is a function of x?

OpenStudy (anonymous):

\[ y= \pm \sqrt{ 1 - x^2} \] So what can you say?

OpenStudy (anonymous):

y is a function of x y is not a function of x

OpenStudy (anonymous):

Is y a function of x. Why or why not?

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

I have to choose either y is a function of x or y is not a function of x

OpenStudy (anonymous):

It is not. Why?

OpenStudy (anonymous):

So why is not a function of x.

OpenStudy (anonymous):

For each value of x between -1 and 1, you get two values of y. You should get one and only one to be a function.

OpenStudy (anonymous):

So your saying why is not a function of x

OpenStudy (anonymous):

y is not a function of x

OpenStudy (anonymous):

it is only a function if x gives one and only one y

OpenStudy (anonymous):

yes. If x =0, then \[ y=\pm 1 \]

OpenStudy (anonymous):

Ok, I understand now.